ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 118, 2000

Arnon, D. Generalized Dickson invariants  183

Briskin, M., J.-P. Francoise and Y. Yomdin Center conditions II: Parametric and model center problems

61

Briskin, M., J.-P. Francoise and Y. Yomdin Center conditions III: Parametric and model center problems

83

Dal’bo, F., J.-P. Otal et M. Peigné Séries de Poincaré des groupes géométriquement finis

109

De Carli, L. Unique continuation for elliptic operators with nonmultiple characteristics

15

Francoise, J.-P. See M. Briskin

Gekeler, E.-U. A note on the finiteness of certain cuspidal divisor class groups

357

Harbater, D. Embedding problems with local conditions  317

Heo, J. Representations of invariant multilinear maps on Hilbert

C^{∗}-modules 125

Kim, H. H. Correction to “Langlands–Shahidi method and poles of automorphic L-functions II”, Israel Journal of Mathematics, Vol. 117, 2000, pp. 261–284

379

Koszmider, P. On strong chains of uncountable functions  289

Lindenstrauss, J., E. Matoušková and D. Preiss Lipschitz image of a measure-null set can have a null complement

207

Matoušková, E. See J. Lindenstrauss

Merzbach, E. and D. Saada Stochastic integration for set-indexed processes

157

Otal, J.-P. See F. Dal’bo

Peigné, M. See F. Dal’bo

Pemantle, R. and Y. Peres Nonamenable products are not treeable

147

Peres, Y. See R. Pemantle

Preiss, D. See J. Lindenstrauss

Rugeihyamu, S. E. On the Koebe and Bieberbach constants for a hyperbolic planar domain 369 Saada, D. See E. Merzbach Schramm, O. Scaling limits of loop-erased random walks and uniform

spanning trees

221

Tambara, D. Representations of tensor categories with fusion rules of self-duality for abelian groups

29

Videla, C. R. Definability of the ring of integers in pro-p Galois extensions of number fields

1

Yomdin, Y. See M. Briskin